# Toward a Physical Theory of Intelligence

### v1

@@ 3.3 Intelligence
Given the goal potential $G:X\times E\to\mathbb{R}$ introduced in §2.2, we define intelligence as a measure of the efficiency with which irreversible information processing is converted into goal-directed physical work. Over a finite horizon $T$, the expected goal-directed work performed by the coupled agent-environment process $\mu_{0:T}$ is
$W_{\mathrm{goal},T}=\mathbb{E}_{\mu_{0}}[G(X_{0},E_{0})]-\mathbb{E}_{\mu_{T}}[G(X_{T},E_{T})],$ (35)
corresponding to the net reduction of the goal potential over the interval.
@@
Let $I_{\mathrm{irr},T}$ denote the total amount of information irreversibly processed over the same horizon,
$I_{\mathrm{irr},T}=\int_{0}^{T}\dot{I}_{\mathrm{irr}}(t)\,dt,$ (36)
@@
where $\dot{I}_{\mathrm{irr}}(t)$ is the instantaneous irreversible information rate defined in §2.3. The cumulative intelligence over $[0,T]$ is then defined as
$\chi_{T}=\frac{W_{\mathrm{goal},T}}{I_{\mathrm{irr},T}},$ (37)
@@
which quantifies the amount of goal-directed work extracted per nat of irreversibly processed information. This ratio is well defined whenever $I_{\mathrm{irr},T}>0$ and characterizes the fundamental efficiency with which irreversible computation is expressed as useful work, independent of the absolute timescale of the underlying dynamics.
@@
For analyses conducted at the level of instantaneous fluxes, define the goal-directed work rate
$\dot{W}_{\mathrm{goal}}(t):=-\,\frac{d}{dt}\,\mathbb{E}_{\mu_{t}}[G(X_{t},E_{t})].$ (38)
The corresponding instantaneous intelligence is
$\chi(t)=\frac{\dot{W}_{\mathrm{goal}}(t)}{\dot{I}_{\mathrm{irr}}(t)},$ (39)
which serves as a local diagnostic of how irreversible information processing is expressed as work at a given moment in time. In practice, intelligence is most naturally evaluated over horizons for which irreversible information processing produces measurable changes in the goal potential, as captured by the cumulative quantity $\chi_{T}$.
The definition of intelligence is agnostic to the specific control, inference, or decision mechanisms employed by the agent. It depends only on the physical evolution of the joint measure $\mu_{t}$ and on the irreversible information transformations generated by that evolution.
3.4 Physical Constraints on Achievable Intelligence
Having defined information capacity, identifiability capacity, and intelligence, we now characterise how these quantities jointly constrain achievable goal-directed behaviour in physical systems. Some of these constraints take the form of explicit inequalities, while others arise as feasibility limits imposed by sensing, inference, and conservation-law–constrained dissipation. Together, they delineate the physically admissible region of performance, independent of the agent’s internal architecture or control strategy.
3.4.1 Information capacity is bounded above by input identifiability.
Let $Z_{t}:=\Gamma_{E}(E_{t})$ and $Y_{t}:=\Gamma_{X}(X_{t})$ denote the environmental input and system output ports. Define an independent copy $Z_{t}^{\prime}$ of $Z_{t}$ (conditioned on $Z_{[0,t)},Y_{[0,t)}$), and let $\overline{\mathcal{F}}_{t}(Z_{t},Z_{t}^{\prime})$ denote the path-averaged Fisher information of the observation channel along the interpolation path between $Z_{t}^{\prime}$ and $Z_{t}$. Define the coupled spread–curvature matrix
$\mathcal{G}_{t}:=\mathbb{E}\!\left[(Z_{t}-Z_{t}^{\prime})(Z_{t}-Z_{t}^{\prime})^{\top}\,\overline{\mathcal{F}}_{t}(Z_{t},Z_{t}^{\prime})\;\middle|\;Z_{[0,t)},\,Y_{[0,t)}\right].$ (40)
Then for any finite horizon $T$,
$\mathcal{C}_{T}[\bar{\mu}_{0:T}]\;\leq\;\frac{1}{T}\int_{0}^{T}\tfrac{1}{2}\,\mathrm{tr}\!\big(\mathcal{G}_{t}\big)\,dt.$ (41)
Thus the rate at which environmental structure can enter the system is bounded by the joint interaction between the conditional spread of inputs and the local curvature (sensitivity) of the observation channel. Intake is maximized along directions in which variations in the input induce large changes in the output distribution. A derivation of (41) is provided in Appendix B.
3.4.2 Information capacity constrains achievable goal-directed work.
Information capacity characterises the rate at which new information about the environment becomes available to the agent through the agent–environment interface. Although information capacity does not directly bound the efficiency with which irreversible information processing is converted into work, it constrains which agent–environment interaction trajectories are informationally admissible over a finite horizon.
Fix a task specification, environment dynamics, and interaction horizon $[0,T]$. Let $\Pi(\mathcal{C}_{T},\mathcal{F}_{T})$ denote the set of agent policies whose induced trajectories are consistent with an interface information capacity $\mathcal{C}_{T}$ and an identifiability capacity $\mathcal{F}_{T}$ along the resulting trajectory. Each policy $\pi\in\Pi(\mathcal{C}_{T},\mathcal{F}_{T})$ induces a corresponding amount of goal-directed work $W_{\mathrm{goal},T}(\pi)$.
The set of achievable goal-directed work values under these informational constraints is therefore

### v2

@@ 3.1 Intelligence
Given the macroscopic causal footprint developed in §2.2, intelligence is primarily evaluated with respect to the expected causal work $W_{\mathrm{causal},T}$ performed on the environment. Over a finite horizon $T$, this is quantified by the expected net extraction of structural capacity over the interval:
$W_{\mathrm{causal},T}=\mathbb{E}_{\mu_{0:T}}[W_{\mathrm{port},T}-\Delta A_{\mathrm{env}}^{\mathrm{causal}}]$ (33)
@@
Let $I_{\mathrm{irr},T}$ denote the total amount of information irreversibly processed by the agent over the same horizon:
$I_{\mathrm{irr},T}=\int_{0}^{T}\dot{I}_{\mathrm{irr}}(t)\,dt$ (34)
@@
where $\dot{I}_{\mathrm{irr}}(t)$ is the instantaneous irreversible information rate defined in §2.3. The cumulative intelligence over $[0,T]$ is then defined as:
$\chi_{T}=\frac{W_{\mathrm{causal},T}}{I_{\mathrm{irr},T}}$ (35)
@@
which quantifies the amount of causal structural work extracted per nat of irreversibly processed information. This ratio is well defined whenever $I_{\mathrm{irr},T}>0$ and characterises the fundamental physical efficiency with which irreversible computation is expressed as macroscopic causal influence, independent of the absolute timescale of the underlying dynamics.
@@
For analyses conducted at the level of instantaneous fluxes, the causal work rate is defined via the expected instantaneous power:
$\dot{W}_{\mathrm{causal}}(t):=\mathbb{E}_{\mu_{t}}[\dot{W}_{\mathrm{port}}(t)-\dot{A}_{\mathrm{env}}^{\mathrm{causal}}(t)]$ (36)
The corresponding instantaneous intelligence is:
$\chi(t)=\frac{\dot{W}_{\mathrm{causal}}(t)}{\dot{I}_{\mathrm{irr}}(t)}$ (37)
which serves as a local diagnostic of how irreversible information processing is expressed as causal influence at a given moment in time.
While raw causal intelligence $\chi_{T}$ measures the absolute scale of an agent’s structural influence, evaluating intelligence in the context of a specific task requires isolating the goal-aligned component of this footprint. To account for this, a projection operator $\pi_{\mathcal{G}}$ that extracts the goal-relevant component of the interface-level footprint can be leveraged. As introduced in §2.2, we define the expected goal-directed work as: